OpenStudy (anonymous):

Find the three arithmetic means between 10 and 18.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what definition is your teacher using?

OpenStudy (anonymous):

find three values that split the numbers equally? 12,14,16?

OpenStudy (anonymous):

The terms between two nonconsecutive terms of an arithmetic sequence

OpenStudy (anonymous):

ok so you want the 3 terms in the middle so that there is an equal distance between all numbers

OpenStudy (anonymous):

(18-10)=8, you want 3+1 equal parts. 8/(3+1)=8/4=2 10+2=12 12+2=14 14+2=16 16+2=18 12,14,16

OpenStudy (anonymous):

i am glad you can make head or tails out of this.